A metallurgist has one alloy containing 29% aluminum and another containing 51% aluminum. How many pounds of each alloy must he use to make 48 pounds of a third alloy containing 50% aluminum?
Im stuck after
x+y=48
.29x+.51y=?

Answer by TimothyLamb(4379)

x = pounds of 29% alloy
y = pounds of 51% alloy
---
x + y = 48
29x + 51y = 50*48
---
put the system of linear equations into standard form
---
x + y = 48
29x + 51y = 2400
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x = pounds of 29% alloy = 2.1818182
y = pounds of 51% alloy = 45.818182
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php